Determine the specific heat of a 335-gram alloy which is at the boiling point of water, and when put into 300 ml of 10°C H2O reaches equilibrium at 70°C.  How much heat is absorbed by this system? (cwater=1 cal/groC)




1
Expert's answer
2021-02-16T10:11:26-0500

t1(alloy) = 100 oC

t2(alloy) = 70 oC

"\Delta"t(alloy) = 100 - 70 = 30 oC

Q = "\Delta"t(water) "\times" m(water)"\times"C(water) = (70 - 10)"\times" 300"\times"= 18000 (cal)

C (alloy) ="\frac{Q}{(\Delta t(alloy) \times m(alloy))}" "\frac{Q}{(\Delta t(alloy) \times m(alloy))}" = "\frac{18000}{(30 \times 335)}" = 1.791(cal "\times" g-1"\times" oC-1)

Answer: C(alloy) = 1.791 (cal "\times" g-1"\times" oC-1)